# Trajectory Optimization for In-Hand Manipulation with Tactile Force Control

### v1

@@ IV-C Open-Loop and Force Controller
Fig. 8 shows the changes in force magnitude from the MTSs on the first finger. Unlike the open-loop controller, where the force magnitude continuously decreases throughout the trajectory, integrating the force controller with the reference trajectory enables the force magnitude to remain close to the desired force magnitude. Although force magnitude tends to oscillate due to the limited controllability of the fingers, the force controller can be stopped at a high force measurement, enabling robust execution of the next task. Since the force controller maintains a certain level of contact force, the state estimator becomes more stable compared to the open-loop case, increasing the likelihood of satisfying the terminal state conditions for the object. Rolling was successfully achieved in 7 out of 10 trials, while the remaining 3 trials exhibited deviations from the ground truth.
@@ V Conclusion
This paper introduces a framework for in-hand manipulation with human-like robotic hands that are equipped with tactile sensors. We present the finger force controller with offline motion planning using a state estimator with tactile sensor and NLP. By demonstrating the rolling of an object with a radius of 7.5 mm using two fingers, we showed that the proposed method can perform robust and accurate rolling motion. As we use relatively compact tactile sensor compared to vision based tactile sensors, this research can be applied to tasks that require precise and robust manipulation of small objects in tight spaces. For future research, we plan to investigate the rolling of objects with complex shapes or deformable objects, where the dynamics themselves are uncertain.
@@ Acknowledgement
This research was supported by Pioneer Center for Accelerating P2X Materials Discovery (CAPeX), DNRF grant number P3

### v2

@@ IV-C Open-Loop and Force Controller
Fig. 8 shows the changes in force magnitude from the MTSs on the first finger. Unlike the open-loop controller, where the force magnitude continuously decreases throughout the trajectory, integrating the force controller with the reference trajectory enables the force magnitude to remain close to the desired force magnitude. Although the force magnitude tends to oscillate. Since the force controller maintains a certain level of contact force, the state estimator becomes more stable compared to the open-loop case, increasing the likelihood of satisfying the terminal state conditions for the object. Rolling was successfully achieved in 7 out of 10 trials, while the remaining 3 trials exhibited deviations from the ground truth.
@@ V Conclusion
This paper introduces a framework for in-hand manipulation with human-like robotic hands that are equipped with tactile sensors. We present the finger force controller with offline motion planning using a state estimator with tactile sensor and NLP. By demonstrating the rolling of an object with a radius of 7.5 mm using two fingers, we showed that the proposed method can perform robust and accurate rolling motion. As we use relatively compact tactile sensor compared to vision based tactile sensors, this research can be applied to tasks that require precise and robust manipulation of small objects in tight spaces. For future research, we plan to investigate the rolling of objects with complex or deformable shapes, where the dynamics are uncertain, and aim to achieve a higher success rate.
@@ Acknowledgement
This research was supported by Pioneer Center for Accelerating P2X Materials Discovery (CAPeX), DNRF grant number P3 and Fabrikant Vilhelm Pedersen og Hustrus Legat.